OpenStudy (anonymous):

If h(x) is the inverse of f(x), what is the value of h(f(x))? 0 1 x f(x) Would it be x?

OpenStudy (anonymous):

yes! it would be x

OpenStudy (anonymous):

for example, f(x) = x^2 and h(x) = sqrt(x), then h(f(x)) = sqrt(x^2) = x

OpenStudy (anonymous):

another example, f(x) = e^x and h(x) = ln(x), then h(f(x)) = ln(e^x) = x

OpenStudy (anonymous):

thank you for the explanation!

OpenStudy (anonymous):

welcome!